OpenStudy (anonymous):

solve each system by graphing and tell whether it is consistent or inconsistent if the system is consistent tell whether it is dependent or independent 3x+y=5 y=3x-4

OpenStudy (saifoo.khan):

My fav link, http://www.jcoffman.com/Algebra2/ch3_1.htm

OpenStudy (anonymous):

the x is the one u plot right then u go up usin the y?

OpenStudy (anonymous):

how would you graph x = 5/3 and y = -7/4

OpenStudy (ash2326):

rebecca Let's plot the two graphs first. I'll show you:)

OpenStudy (anonymous):

alright thanks :)

OpenStudy (anonymous):

Everything okay. you were typing for a while and suddenly stopped

OpenStudy (ash2326):

First equation is \[3x+y=5\] Let's plot this first put x=0 in this equation , we get y=5 so one point is (0, 5) put y=0, we get 3x=5 or x=5/3 so another point is (5/3, 0) Now we'll mark these two points on the graph and then join them together